OpenStudy (anonymous):

ASAP please Find the area of the region R bounded by x- y^2 = 0 x+7y^2 =8

OpenStudy (abb0t):

find your boundaries first by setting your functions equal to each other and finding "x"

OpenStudy (anonymous):

y= +/- 1

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

x = +- 1, too

OpenStudy (anonymous):

ok, I am studying this stuff, too; we can do together

OpenStudy (anonymous):

sketch the graph first, friend, you do this step, next one will be mine,OK?

OpenStudy (anonymous):

so, let you do alone,

OpenStudy (anonymous):

i don't know how to graph it on this thing

OpenStudy (anonymous):

i've got til 10 to do this problem :/

OpenStudy (anonymous):

so, you just need the answer?

OpenStudy (anonymous):

before ten, and then i need to work it out for my own sake

OpenStudy (anonymous):

ok, I set up the int for you y goes from -1 to 1 , x goes from y^2 to 8 - 7y^2 the function is 1 dxdy you can do it , right?

OpenStudy (anonymous):

in two minutes? i doubt it

OpenStudy (anonymous):

:(

OpenStudy (abb0t):

I think it might be easier if you graphed it as: \(x = y^2\) and \(x =8 -7y^2\) since those are graphs that you can easily graph.

OpenStudy (abb0t):

@Hoa are you using greens theorem to evaluate this?

OpenStudy (anonymous):

no, hihi... just draw the graph, solve for intersection point, take double integral

OpenStudy (anonymous):

what is th greens theorem?

OpenStudy (abb0t):

I think you can solve this using single integrals since you're using area "R".

OpenStudy (anonymous):

how?show me ,please. I need it for my final